BexleyBrooks:

Need some more help on Math. 1. x-9 over 5=4 what is x? 2. a+12 over 2=-8 what is a? 3. 7=c-16 over 8 what is c? 4. -3=4+r over 6 what is r? 5. 13-x=7 over 3 what is x? 6. x+3(x+4)=20 what is x? 7. 6(y-1)+8=32 what is y?

GIGI18:

idk T-T

GIGI18:

what grade math is this

luhivqqcherry:

7. 6(y-1)+8=32 what is y? y = 5 7 explantion= 6(y−1)+8−8=32−8 6(y-1) + 0 = 24 6(y−1)=24

luhivqqcherry:

y−1+1=4+1 y−0=5y=5

luhivqqcherry:

@luhivqqcherry wrote:
y−1+1=4+1 y−0=5y=5
this one is from number 7

jhonyy9:

1. \[\frac{ x-9 }{ 5 } = 4 \] multiplie both sides by 5 so will get x-9 = 20 add to both sides 9 x = 29 in the same way try the next too

HotPockets14:

1.)x = 29 2.) a = -28 3) c = 9 4) r = -42 5) x = 32/3 6) x = 2 7) y = 5

BexleyBrooks:

@gigi18 wrote:
what grade math is this
8th

jhonyy9:

@hotpockets14 wrote:
1.)x = 29 2.) a = -28 3) c = 9 4) r = -42 5) x = 32/3 6) x = 2 7) y = 5
step by step please ,dont give dirrect answers never